OpenStudy (anonymous):

A quantity of gas with an initial volume of 1 cubic foot and a pressure of 500 psi expands to a volume of 2 cubic feet. Find the work done by the gas...Hint: Find K First...Please help!

OpenStudy (anonymous):

Do you know the formula?

OpenStudy (anonymous):

Do you use something like: \[ PV = kT \]

OpenStudy (anonymous):

I know its W= the integral from V1 to V0 of K/V dV

OpenStudy (anonymous):

Or that P=k/V

OpenStudy (anonymous):

Oh, so we can assume temperature does not change?

OpenStudy (anonymous):

I know the pressure varies inversely as volume.

OpenStudy (anonymous):

Okay so it really is just \[ P(V) = \frac{K}{V} \implies K = PV \]And that means work is: \[ W = \int_{V_1}^{V_2} P(V) dV = \int_{V_1}^{V_2} \frac{K}{V} dV = K\left[\ln(V)\bigg|_{V_1}^{V_2}\right] = K\ln\left(\frac{V_2}{V_1}\right) = P_1V_1\ln\left(\frac{V_2}{V_1}\right) \]

OpenStudy (anonymous):

We could technically use any pair of \(P\) and \(V\) at any moment, but since we are given the initial pressure and volume, we'll use those.

OpenStudy (anonymous):

Okay so I will just replace P1 with 500 psi and V1 with 2ft^3?

OpenStudy (anonymous):

Yeah...

OpenStudy (anonymous):

Okay but that will give me some really crazy answer! Are you sure?

OpenStudy (anonymous):

Umm, you have to be careful about your units.

OpenStudy (anonymous):

Okay well what do you think the best units to use would be?